OpenStudy (anonymous):

You can go 630 miles on 35 gallons of gas. How many miles per gallon does your car go?

thomaster (thomaster):

35 gallon = 630 miles 1 gallon = ?? miles \(\Large\dfrac{630}{35}= \large\sf miles\ per\ gallon\)

OpenStudy (anonymous):

18

thomaster (thomaster):

yep :)

OpenStudy (anonymous):

wat about The perimeter of a rectangular lot is 56 meters. The length exceeds the width by 12 meters. Find the length of the lot.

thomaster (thomaster):

 width = x length = x+12 So the equation for the perimeter is: \(\large 4x+24=56\) now solve for x and add 12

OpenStudy (anonymous):

48

OpenStudy (anonymous):

?

thomaster (thomaster):

what is the width?

thomaster (thomaster):

Do you know how to solve \(\large 4x+24=56\) for x?